Find a number such that one-half of the number is 5 less than two-thirds of the number.
1
Expert's answer
2018-11-08T14:13:09-0500

Let the number be x. Then:

x/2=(2x/3)-5 <=>

(2x/3)-(x/2)=5 <=>

(4x/6)-(3x/6)=5 <=>

x/6=5 =>

x=30